Lowering Detector

Lowering Detector

Lowering Detector

Total weight 20,000 tons

yb0 1800 tons

ye1 1400 tons

115% test has been run

Magnetic Field

4.0 Tesla

2112 turns 20.0 kA

12.6 H

13m long 3 meter radius

2.7 GJoules


10 million silicon microstrips (10cm x 80 um)

80 million pixels (100x150 um’s) (25x180 um)

sigma_vertex r,phi 10 um z 20 um

200 square meters of silicon (~200 x larger than LEP era)

tracker layout p44

Status: TIB/TID/TOB/TEC modules completed

Module integration well under way.

Transportation to Point 5 Jan07

Pixels also in good shape. Pixel installation

after LHC pilot run. Jan08

June 10th:The final set of CMS tracker outer barrel silicon detectors left Lab B at Fermilab for delivery to the CERN laboratory in Geneva, Switzerland...

Inner Barrel (TIB), Outer Barrel (TOB), Inner Disks (TID) and Endcap (TEC).

In October it is expected that the first TIB half will be inserted into the TOB for the first time with combined commissioning expected to start soon after.

Electromagnetic Calorimeter

80,000 lead tungstate crystals

barrel APD (avalance photo-diodes)

endcap VPT (vacuum photo-triodes)


EB: Crystal production complete March 1,2007. Crystal production on schedule.

Modules complete end of April 2007.

EE: Crystals start arriving Sept 06

Assembly start Mid-March 2007.

DEE1 before pilot run Sept 07

DEE2 Dec 07

DEE3 mid Mar 08

DEE4 mid Apr 08

Hadronic PreShower

Two layers silicon detectors

Factor of 3x rejection of pi0 from gamma trigger

Hadronic Calorimeter

Barrel/Endcap:Brass Tiles, imbedded WLS fibers

Forward: Quartz Crystals imbedded in copper

Hybrid Photodiodes (HPD)

Optical readout ganged into HB1,HB2 for 17 brass/scint layers ad HB3 for outer two scint. layers

Readout 25 nsecs , conditioned ADC 3 bits range, 5 bits charge

ME 4176 channels

MB 7056 channels

MO 3864 channels

Status: HCAL MO,ME,MB done.

Hardware and Electronics Completed.


468 chambers

200 um resolution (phi)

with 6 planes unfolded area of football field

ArCo2CF4 40:50:20

ganged wires (discriminator) 180K

strips (ADC 50nsec sampling) 240K

Status: 92% CSC installed/commissioned

Off chamber electronics 50% installed commissioned

HV production complete, 17% installed

Drift Tubes

250 chambers ArCo2 20%

v(drift)=55um/nsec -> total drift time ~700nsec

200 um resolution (phi)

Amp/Discr/TDC 200K channels

Status: Stats 144 chambers installed.

76 RPC+DT installation Nov 1,06-Dec 22,06.


Coarse position, fine timing

Barrel 80640 channels (amp/discr)

Endcap 92162 channels (amp/discr)

Status: Barrel Chambers completed, Mounting Integration ongoing

Endcap Chambers completed by end of 2006


3.2 usec latency

Regional Trigger Calorimeter (RCT) Global Muon Trigger(GMT) Global Trigger (GT) Global Calorimeter Trigger (GCT)

Status: RPC completion Nov 2006, CSC done, DT completion Nov 2006

ECAL barrel Sept 06, endcap Summer 07

HCAL done

RCT done

GCT completion Mar 07

GT & GMT done


event size 1.5 Mbytes

Farm Reduction 100 KHz->100 Hz

Status: Sept 06 purchase 600 DAQ PC’s

600 DAQ PC’s Dec 06 (50KH rate)

600 more DAQ PC’s Dec 08 (100KHz rate)


Event rate 100 Hz x 2 Meg (event size) = 200 MB/s => 20 TB/day (probably 10 TB/day

with beam spills)

30 tape driver (~1 million $)

750 TB disk

presently used 5.7 FB total (~1 million $ tapes)

results 2.2GB/s =1 GB/s readback + 1.2 GB/s to tape

example IBM3592 40MB/s 600GB/tape $32,000

Early Running at sqrt(s)=900 GeV

Design luminosity 2x10^33

More detector computer commissioning than physics

Few W’s and Z’s

Luminosity ~10^29

Sigma(mb) 50mb (UA5)

Minimum bias rate 5 Hz

Last Magnet Delivered Oct 2006

Last Magnet Tested Mar 2007

Machine Closed Aug 2007

First Collision Nov 2007

¼ magnets commissioned to 7TeV, rest commissioned to 1.2 TeV degaussing limit

rest (3/4) of magnets will be commissioned to 7TeV in winter 2007-8 shutdown


Physics TDR